Lexi Belcher picked up the monthly report that Irvin Santamaria left on her desk. She smiled as her eyes went straight to the bottom line of the report and saw the favorable variance for operating income, confirming her decision to push the workers to get those last 270 cases off the production line before the end of the month.

But as she glanced over the rest of numbers, Lexi couldn’t help but wonder if there were errors in some of the line items. She was puzzled at how most of the operating expenses could be higher than the budget since she had worked hard to manage the production line to improve efficiency and reduce costs. Yet the report, shown below, showed a different story.

Actual Budget Variance
Cases produced and sold 10,250 9,980 270 Favorable
Sales revenue $1,943,600 $1,866,300 $77,300 Favorable
Less variable expenses
   Direct material 559,878 548,900 10,978 Unfavorable
   Direct labor 267,115 259,480 7,635 Unfavorable
   Variable manufacturing overhead 284,442 279,440 5,002 Unfavorable
   Variable selling expenses 92,944 89,820 3,124 Unfavorable
   Variable administrative expenses 41,657 39,920 1,737 Unfavorable
Total variable expense 1,246,036 1,217,560 28,476 Unfavorable
Contribution margin 697,564 648,740 48,824 Favorable
Less fixed expenses
   Fixed manufacturing overhead 110,778 109,780 998 Unfavorable
   Fixed selling expenses 69,361 69,860 (499 Favorable)
   Fixed administrative expenses 129,540 129,740 (200 Favorable)
Total fixed expense 309,679 309,380 299 Unfavorable
Operating income $387,885 $339,360 $48,525 Favorable



Lexi picked up the phone and called Irvin. “Irvin, I don’t get it. We beat the budgeted operating income for the month, but look at all the unfavorable variances on the operating costs. Can you help me understand what’s going on?” “Let me look into it and I’ll get back to you,” Irvin replied.

Irvin gathered the following additional information about the month’s performance.

Direct materials purchased: 101,796 pounds at a total of $559,878
Direct materials used: 101,796 pounds
Direct labor hours worked: 26,447 at a total cost of $267,115
Machine hours used: 40,868



Irvin also found the standard cost card for a case of product.

Standard Price Standard Quantity Standard Cost
Direct materials $5.50 per pound 10 pounds $55
Direct labor $10 per DLH 2.59 DLH 25.90
Variable overhead $7 per MH 4 MH 28.00
Fixed overhead $2.74 per MH 4 MH 10.96
Total standard cost per case $119.86

Calculate the direct material price variance and direct material quantity variance for the month. (If variance is zero, select "Not Applicable" and enter 0 for the amounts.)

Direct material price variance $

Enter dollar amount

Favorable / Unfavorable / Not Applicable

Direct material quantity variance $

Enter Dollar amount

Unfavorable / Not Applicable / Favorable
Answer #1

Direct material price variance = (Standard price-actual price)actual quantity = (5.5*101796-559878) = 0 None

Direct material quantity variance = (Standard qty-actual qty)Standard price = (102500-101796)*5.5 = 3872 F